The Number of Walks on a Regular Tree

Eric ROWLAND^{1} and Doron ZEILBERGER^{1}

1 Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen

Rd., Piscataway, NJ 08854-8019, USA. [erowland,zeilberg] at math dot rutgers dot edu ,
http://www.math.rutgers.edu/~[erowland, zeilberg] . First version: xx.x, 200x.